OpenStudy (anonymous):

What is the slope of the line that passes through the points (2,5) and (7,3)? A. -5/2 B. -2/5 C. 8/9 D. 9/8

OpenStudy (anonymous):

B.

OpenStudy (anonymous):

B. -2/5

OpenStudy (anonymous):

To find the slope of a line given two points, simply divide the change in Y by the change in X: \[\Large \frac{{\Delta Y}}{{\Delta X}}\] So in this case: \[\Large \frac{{3 - 5}}{{7 - 2}} = \frac{{ - 2}}{5}\]